OpenStudy (anonymous):

Find an equation of each line with the given slope that passes through the given point. Write the equation in the form Ax+By=C. Q: m=3/2 ; (5,-6) Need help working it out. I'm not too good with fractions.

OpenStudy (anonymous):

\[y+6=\frac{3}{2}(x-5)\] is a start

OpenStudy (anonymous):

to get rid of the fraction, multiply both sides by \(2\)

OpenStudy (anonymous):

Well what do I do about the 5 at the end ?

OpenStudy (anonymous):

to get rid of thefraction mutlply both sides

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

leave it in the parentheses for the moment

OpenStudy (anonymous):

bye 2 then put them togher

OpenStudy (anonymous):

\[y+6=\frac{3}{2}(x-5)\\ 2y+12=3(x-5)\]

OpenStudy (anonymous):

and put the biggest one on the top

OpenStudy (anonymous):

@satellite73 Thank you again!

OpenStudy (anonymous):

yw